# Symbolic Higher-Order Analysis of Multivariate Time Series

### v1

@@ Introduction.
Complex systems exhibit collective behaviours of extraordinary richness. Notable examples include the emergence of cooperation [1, 2] and epidemics in social groups [3], as well as synchronization in biological systems [4] and in human cognition [5]. Such behaviors directly stem from the interactions between the many fundamental units of such systems, e.g. individuals in a society or neurons within the brain [6, 7, 8, 9]. The most intuitive way to represent interactions in a complex system is through a network whose nodes correspond to the fundamental units of the complex systems and the edges describe direct or indirect couplings, or other dependencies or correlations between two units [10, 11, 12]. However, networks often provide an oversimplified description of a complex system in terms of pairwise relations. In many cases, units interact in groups, and these group interactions cannot be reduced to pairwise relationships among group members [13]. Higher-order (HO) representations of the interaction structure, such as simplicial complexes and hypergraphs, recently have been extensively and successfully applied to the study of complex systems [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. In many cases, the interactions within a complex system must be inferred directly from the observation of the system dynamics, specifically from multivariate time series obtained by measuring the activity of each of the units of the system. Given the importance of the problem, numerous methods have been developed over the years to reconstruct pairwise networks from such empirical time-series data [25, 26, 27, 28, 29]. In contrast, efforts to reconstruct or filter higher-order interactions have only emerged more recently [30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. Existing approaches typically rely on one of the following: the assumption of continuous, differentiable time series [33, 34]; discretized versions of continuous time series with specific statistical properties (e.g., normally distributed values) [35, 37, 38], or prior knowledge of the system’s underlying dynamical process and its functional form [33, 39]. These assumptions are often too restrictive in real-world scenarios, where the exact dynamical rules generating the time series are unknown [34], and the distribution of values across real-world datasets can be highly heterogeneous [40]. Additionally, the activity of individual units is frequently neither continuous nor differentiable; instead, it tends to occur at discrete moments in time, with events that are effectively instantaneous relative to the typical timescale of observation. Relevant examples are the spiking activity of neurons, seismic events, user activity on social platforms and sell-buy orders in stock markets. In all such cases, the dynamics is better described by a set of binary time series data of 00 and $1$, where the $1$ states represent the given points in time at which the highly localized events occur.
@@
Figure 1: (a)  Each time series $x_{i}(t)$, with $i=1,\ldots N$, is associated to one of $N$ different symbols/colours. (b) Then the set of time series is converted into a symbolic sequence: if an event $x_{i}(t)=1$ is followed by $x_{j}(t)=1$ within a time interval $\Delta t$ (grey bar in (a)), their corresponding symbols are placed adjacently in the symbolic sequence. Otherwise (pink cross) a “space” symbol is inserted in the sequence. (c-d) Lists of ordered and unordered $2$-tuples and $3$-tuples found in the sequence. (e) Statistically significant tuples (motifs) are finally the hyperedges of an hypergraph.
@@
Let us consider a multivariate time series of $N$ binary signals $x_{i}(t)\in\{0,1\},\forall i=1,...,N$, with $t\in[0,T]$, where $x_{i}(t)=1$ represents the occurrence of event of type $i$ at time $t$, and $T$ is the observation time. We want to construct a hypergraph whose $N$ nodes are associated with the different types of events, and the hyperedges represent groups of correlated events. In order to do this, we first convert the multivariate time series into a symbolic sequence, Fig. 1(a,b). We consider an alphabet $\cal A$ of $N+1$ symbols, the first $N$ symbols (colours in Figure) respectively associated to each one of the $N$ time series, plus a special “space” (or empty) symbol. We then construct an ordered sequence $\cal S$ of such symbols from the original set of $N$ time series in the following way. When the event $x_{i}(t)=1$ is followed by the event $x_{j}(t^{\prime})=1$ within a given time interval $\Delta t$ (i.e., when $t^{\prime}-t<\Delta t$), the symbols corresponding to $i$ and $j$ are placed in adjacent positions in the symbolic sequence $\cal S$. If instead no activity occurs within $\Delta t$, the space symbol is inserted in the sequence after $i$. Finally, from the symbolic sequence $\cal{S}$, we extract the multiset (i.e., with repetitions of elements) $\mathcal{D}^{l}$ of all overlapping $l$-tuples, with cardinality $|\mathcal{D}^{l}|=M^{l}$, and the set $\mathcal{T}^{l}$ of the unique $l$-tuples (without repetition), with cardinality $|\mathcal{T}^{l}|=N^{l}$, see Fig. 1(c,d). A $l$-tuple $\bm{s}=(s_{1},s_{2},\dots,s_{l})\in\mathcal{T}^{l}$ is an ordered (or unordered) sequence of $l$ symbols $s_{i}\in\mathcal{A}~{}\forall i=1,2,\ldots l$. In this way, from a sequence $\mathcal{S}$ of length $S$, we get $M^{l}=S-l+1$ tuples of length $l$.
This approach allows to build a hypergraph $\mathcal{H}(\mathcal{N},\mathcal{E})$ of higher-order correlations in the original multivariate time-series. The $N$ nodes of the hypergraph correspond to the $N$ units (time series), i.e. the different types of possible events. The hyperedges of different sizes $l$, with $l\geq 2$, representing groups of $l$ units whose dynamics are correlated, are defined by the $l$-motifs in $\cal S$, i.e. by the statistically significant $l$-tuples of symbols. This is shown in Fig. 1(e) for the case of $l=2$ and $3$. In order to assess the significance of a $l$-tuple $\bm{s}$, we evaluate the expected probability $p_{\text{exp}}(\bm{s})$ from the observed occurrences of tuples of shorter length:
$p_{\text{exp}}(\bm{s}=s_{1},s_{2})$ $=p_{\text{obs}}(s_{1})p_{\text{obs}}(s_{2})$ (1)
@@
$p_{\text{exp}}(\bm{s}=s_{1},\dots,s_{l})$ $=\frac{p_{\text{obs}}(s_{1},\dots,s_{l-1})p_{\text{obs}}(s_{2},\dots,s_{l})}{p_{\text{obs}}(s_{2},\dots,s_{l-1})}$ (2)
for $l>2$. The observed probability of $\bm{s}$ in the equations above is computed as:
$p_{\text{obs}}(\bm{s})=\frac{n_{\text{obs}}(\bm{s})}{\sum_{s\in\mathcal{T}^{l}}n_{\text{obs}}(\bm{s})}=\frac{n_{\text{obs}}(\bm{s})}{S-l+1}$ (3)
where $n_{\text{obs}}(\bm{s})$ counts the number of occurrences of $\bm{s}$ in $\mathcal{D}^{l}$ and therefore satisfies $\sum_{s\in\mathcal{T}^{l}}n_{\text{obs}}(\bm{s})=S-l+1$. In this way, when evaluating $p_{\text{exp}}(\bm{s})$, we take into account the effect of lower-order correlations coming from tuples of length smaller than $l$ [43]. If we are interested in the case of unordered tuples, we can simply sum over $S_{p}(\bm{s})$, the set of unique permutations $\bm{\sigma}$ of the ordered tuple $\bm{s}=s_{1},\dots,s_{l}$, obtaining: $p_{\text{exp}}^{\text{und}}(\bm{s})=\sum_{\bm{\sigma}\in S_{p}(\bm{s})}p_{\text{exp}}(\bm{\sigma})$ and $n_{\text{obs}}^{\text{und}}(\bm{s})=\sum_{\bm{\sigma}\in S_{p}(\bm{s})}n_{\text{obs}}(\bm{\sigma})$, where $p_{\text{exp}}(\bm{\sigma})$ and $n_{\text{obs}}(\bm{\sigma})$ are respectively the expected probability and the observed occurrences of the permutation $\bm{\sigma}$ for the ordered case.
We then use a Bayesian approach to reject or accept the null hypothesis that each $l$-tuple $\bm{s}_{i}\in\mathcal{T}^{l}$ where $i\in 1,\dots,N^{l}$ appears with probability $p_{\text{exp}}(\bm{s}_{i})$. Let us define the vector of the observed count of $l$-tuples in $\mathcal{S}$: $\mathbf{n}^{l}_{\text{obs}}=\left[n_{\text{obs}}(\bm{s}_{1}),\dots,n_{\text{obs}}(\bm{s}_{N^{l}})\right]$, where $n_{\text{obs}}(\bm{s}_{i})$ is the number of times each unique tuple $\bm{s}_{i}\in\mathcal{T}^{l}$ is repeated in $\mathcal{D}^{l}$. We assume that each $l$-tuple represents the outcome of a random experiment over $S-l+1$ independent trials. That is, the outcome of each trial can be tuple $\bm{s}_{i}\in\mathcal{T}^{l}$ with probability $p(\bm{s}_{i})$, such that $\sum_{i}p(\bm{s}_{i})=1$. Under this assumption, the likelihood of $\mathbf{n}^{l}_{\text{obs}}=\left[n_{\text{obs}}(\bm{s}_{1}),...,n_{\text{obs}}(\bm{s}_{N^{l}})\right]$ is a multinomial distribution:
$\mathcal{L}(\mathbf{n}_{\text{obs}}^{l}\mid\mathbf{p}^{l})=\frac{(S-l+1)!}{\prod_{i=1}^{N^{l}}n_{\text{obs}}(\bm{s}_{i})!}\prod_{i=1}^{N^{l}}p(\bm{s}_{i})^{n_{\text{obs}}(\bm{s}_{i})}$ (4)
where $\mathbf{p}^{l}=\left[p(\bm{s}_{1}),\dots,p(\bm{s}_{N^{l}})\right]$ is the vector of the occurrence probability of each tuple. For the prior $\Pi(\mathbf{p}^{l})$, as commonly done, we take the conjugate distribution of the multinomial distribution, that is the Dirichlet distribution:
@@
where $\alpha_{i}^{l}$ is the concentration parameter of the tuple $\bm{s}_{i}$, which can be interpreted as prior counts or pseudo-counts, and $\Gamma(\cdot)$ is the Gamma function. We adopt an empirical Bayes approach, where part of the information of the data is used to compute the prior. Namely, to compute the concentration parameter of an $l$-tuple, we use the expected probability in Eqs. (1),(2), functions of the observed probability of $l-1$ and $l-2$-tuples: $\alpha_{i}^{l}=n_{\text{exp}}(\bm{s}_{i})+\epsilon$. Here, $n_{\text{exp}}(\bm{s}_{i})=p_{\text{exp}}(\bm{s}_{i})(S-l+1)$ is the expected number of occurrences of the $l$-tuple $\bm{s}_{i}$ in the sequence, and $\epsilon\geq 0$ is a regularization parameter (a minimum number of pseudo-counts) that ensures that the prior is not too confident, in particular for very small $n_{\text{exp}}(\bm{s}_{i})$, and allows for Bayesian updating. For our results we choose $\epsilon=1$. We then use the likelihood to update the prior obtaining the posterior $P(p\mid\text{Data})$ which, given that the multinomial and the Dirichlet distributions are conjugate, is also a Dirichlet distribution:
$P(\mathbf{p}\mid\text{Data})=\text{Dirichlet}($ $\alpha_{1}^{l}+n_{\text{obs}}(\bm{s}_{1}),\alpha_{2}^{l}+n_{\text{obs}}(\bm{s}_{2}),$
$\dots,\alpha_{N^{l}}^{l}+n_{\text{obs}}(\bm{s}_{N^{l}}))$ (6)
@@
To assess the significance of the $l$-tuple $\bm{s}_{i}$, we compute the Jensen-Shannon distance $d_{\text{JS}}(\bm{s}_{i})$ between the marginal distribution of the prior and that of the posterior[44, 45, 46]. Notice that, the marginal of the Dirichlet distribution $(p(\bm{s}_{1}),\dots,p(\bm{s}_{N}))\sim\text{Dirichlet}(\alpha_{1},\alpha_{2},\dots,\alpha_{N})$ is the Beta distribution:
$p(\bm{s}_{i})\sim\text{Beta}(\alpha_{i},\alpha_{0}-\alpha_{i})$ (7)
@@
where $\alpha_{0}=\sum_{i}\alpha_{i}$. If $d_{\text{JS}}(\bm{s}_{i})$ is large, the null hypothesis that the probability of observing $\bm{s}_{i}$ can be estimated from the probabilities of the shorter length tuples in $\bm{s}_{i}$ must be rejected. Hence, the $l$-tuple shows pure correlations of order $l$, is considered a $l$-motif and, as such, is an $l$-hyperedge (a hyperedge with $l$ nodes) of the hypergraph $\mathcal{H}(\mathcal{N},\mathcal{E})$ of higher-order relations in the original time series. Our method naturally allows to associate weights in $[0,1]$ to the hyperedges, as the JS distance ranges between 00 and $1$. Being $d_{\text{JS}}$ a metric, it does not take into account for the relative position of the prior and posterior distributions. We therefore associate to the hyperedge $\bm{s}_{i}$ a weight equal to $d_{\text{JS}}(\bm{s}_{i})$ only when the posterior is concentrated around larger values than the prior. That is, we consider as potentially significant only those motifs for which the mean of the prior is smaller than the mean of posterior, but a similar analysis could be done also for under-represented motifs. An unweighted version of the hypergraph $\mathcal{H}$ can then be obtained by retaining only $l$-hyperedges whose weights are larger than a given threshold $d_{\text{JS}}^{\text{thr}}(l)$ (see SM for a discussion on how to choose such a threshold). In this way, the Jensen-Shannon distance between the prior and a posterior distributions of each tuple becomes a score of statistical significance; hereinafter we will indicate it as the BJS-score (Bayesian-Jensen-Shannon).
@@
Fig. 2 shows the Receiver Operating Characteristic (ROC) and Precision-Recall (PR) curves illustrating the performance of the BJS-score and the $z$-score in detecting motifs, for noise distributed according to Zipf’s law with exponent $\gamma=1$. See the Supplemental Material (SM) for similar results for other values of $\gamma$. From the left to the right, the curve points correspond to values of decreasing inference thresholds (from $1$ to 00 for the BJS, from $\infty$ to 00 for the z-score). The quicker the ROC curve rises to the upper left corner, the better the inference. On the other hand, the slower the Precision-Recall curve decreases to the bottom right corner, the better. We are interested in particular in having a high precision with few false positive. Our dataset is unbalanced, with many true negative (and the number of true negatives increases as the length of the sequence, i.e. the $r_{\text{ns}}$ increases) therefore the Precision-Recall curve is more informative. The results show that our method outperforms the $z$-score in all scenarios and is capable of detecting motifs even in extreme cases with a noise-to-signal ratio of $r_{ns}=100$. Fig. 2 shows the ROC and PR curves for the combined recognition of $2$- and $3$-motifs. For a separate analysis of $2$- and $3$-motifs, see the Supplemental Material (SM). Moreover, we found that for all the values of $\gamma$ explored, the optimal threshold for the BJS-score, corresponding to the maximum value of F1 score, is $\text{BJS}^{\text{thr}}(l)\simeq[0.5,0.7]$ for both $l=2$ ($2$-tuples) and $l=3$ ($3$-tuples). $\text{BJS}^{\text{thr}}\geq 0.5$ has also an intuitive interpretation; as the Jensen-Shannon distance takes value in $[0,1]$, where 00 corresponds to identical distributions and $1$ to completely different ones. Hence a distance greater than the middle point $0.5$ can be intuitively seen as two distributions that are more different than similar. In the next session on applications to real-world datasets, we will adopt a single threshold $\text{BJS}^{\text{thr}}=0.6$ for both $l=2$ and $l=3$.
@@
For the application to the brain, we studied the spiking activity at the level of individual neurons in the mouse during visual tasks. The multivariate time-series were obtained from the electrophysiological recordings in the Neuropixel datasets of the Allen Brain Observatory [48]. We set $x_{i}(t)=1$ to indicate that an action potential occurred in the neuron $i$ at time $t$. We considered two different spatial scales, one that involves individual neurons ($N=346$), and the other that aggregates neurons in the same functional area ($N=15$, see SM). Higher-order motifs were present in both situations. However, only at the macro-scale of functional brain areas the number $E_{3}$ of $3$-motifs exceeds the number $E_{2}$ of pairwise interactions (more than $70\%$ of the total number of hyperedges). These results are in agreement with a recent study of large-scale electroencephalographic (EEG) brain signals [34], where a similar percentage of brain dynamics was considered HO. The fact that the presence of 3 motifs significantly increases when moving from the microscale of the single neurons to the macroscale of the functional brain areas, suggesting a non-negligible spatial effect that should be taken into account in future studies.

### v2

@@ Introduction.
Complex systems exhibit collective behaviours of extraordinary richness. Notable examples include the emergence of cooperation [1, 2] and epidemics in social groups [3], as well as synchronization in biological systems [4] and in human cognition [5]. Such behaviors directly stem from the interactions between the many fundamental units of such systems, e.g. individuals in a society or neurons within the brain [6, 7, 8, 9]. The most intuitive way to represent interactions in a complex system is through a network whose nodes correspond to the fundamental units of the complex systems and the edges describe direct or indirect couplings, or other dependencies or correlations between two units [10, 11, 12]. However, networks often provide an oversimplified description of a complex system in terms of pairwise relations. In many cases, units interact in groups, and these group interactions cannot be reduced to pairwise relationships among group members [13]. Higher-order (HO) representations of the interaction structure, such as simplicial complexes and hypergraphs, recently have been extensively and successfully applied to the study of complex systems [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. In many cases, the interactions within a complex system must be inferred directly from the observation of the system dynamics, specifically from multivariate time series obtained by measuring the activity of each of the units of the system. Given the importance of the problem, numerous methods have been developed over the years to reconstruct pairwise networks from such empirical time-series data [25, 26, 27, 28, 29]. In contrast, efforts to reconstruct or filter higher-order interactions have only emerged more recently [30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. Existing approaches typically rely on one of the following: the assumption of continuous, differentiable time series [33, 34]; discretized versions of continuous time series with specific statistical properties (e.g., normally distributed values) [35, 37, 38], or prior knowledge of the system’s underlying dynamical process and its functional form [33, 39]. These assumptions are often too restrictive in real-world scenarios, where the exact dynamical rules generating the time series are unknown [34], and the distribution of values across real-world datasets can be highly heterogeneous [40]. Additionally, the activity of individual units is frequently neither continuous nor differentiable; instead, it tends to occur at discrete moments in time, with events that are effectively instantaneous relative to the typical timescale of observation. Relevant examples are the spiking activity of neurons, seismic events, user activity on social platforms and sell-buy orders in stock markets. In all such cases, the dynamics is better described by a set of binary time series data of 0 and $1$, where the $1$ states represent the given points in time at which the highly localized events occur.
@@
Figure 1: (a)  Each time series $x_{i}(t)$, with $i=1,\ldots N$, is associated to one of $N$ different symbols/colours. (b) Then the set of time series is converted into a symbolic sequence: if an event $x_{i}(t)=1$ is followed by $x_{j}(t)=1$ within a time interval $\Delta t$ (grey bar in (a)), their corresponding symbols are placed adjacently in the symbolic sequence. Otherwise (pink cross) a “space” symbol is inserted in the sequence. (c-d) Lists of ordered and unordered 22-tuples and 33-tuples found in the sequence. (e) Statistically significant tuples (motifs) are finally the hyperedges of an hypergraph.
@@
Let us consider a multivariate time series of $N$ binary signals $x_{i}(t)\in\{0,1\},\forall i=1,...,N$, with $t\in[0,T]$, where $x_{i}(t)=1$ represents the occurrence of event of type $i$ at time $t$, and $T$ is the observation time. We want to construct a hypergraph whose $N$ nodes are associated with the different types of events, and the hyperedges represent groups of correlated events. In order to do this, we first convert the multivariate time series into a symbolic sequence, Fig. 1(a,b). We consider an alphabet $\cal A$ of $N+1$ symbols, the first $N$ symbols (colours in Figure) respectively associated to each one of the $N$ time series, plus a special “space” (or empty) symbol. We then construct an ordered sequence $\cal S$ of such symbols from the original set of $N$ time series in the following way. When the event $x_{i}(t)=1$ is followed by the event $x_{j}(t^{\prime})=1$ within a given time interval $\Delta t$ (i.e., when $t^{\prime}-t<\Delta t$), the symbols corresponding to $i$ and $j$ are placed in adjacent positions in the symbolic sequence $\cal S$. If instead no activity occurs within $\Delta t$, the space symbol is inserted in the sequence after $i$. Finally, from the symbolic sequence $\cal{S}$, we extract the multiset (i.e., with repetitions of elements) $\mathcal{D}^{l}$ of all overlapping $l$-tuples, with cardinality $|\mathcal{D}^{l}|=M^{l}$, and the set $\mathcal{T}^{l}$ of the unique $l$-tuples (without repetition), with cardinality $|\mathcal{T}^{l}|=N^{l}$, see Fig. 1(c,d). A $l$-tuple $\boldsymbol{s}=(s_{1},s_{2},\dots,s_{l})\in\mathcal{T}^{l}$ is an ordered (or unordered) sequence of $l$ symbols $s_{i}\in\mathcal{A}~\forall i=1,2,\ldots l$. In this way, from a sequence $\mathcal{S}$ of length $S$, we get $M^{l}=S-l+1$ tuples of length $l$.
This approach allows to build a hypergraph $\mathcal{H}(\mathcal{N},\mathcal{E})$ of higher-order correlations in the original multivariate time-series. The $N$ nodes of the hypergraph correspond to the $N$ units (time series), i.e. the different types of possible events. The hyperedges of different sizes $l$, with $l\geq 2$, representing groups of $l$ units whose dynamics are correlated, are defined by the $l$-motifs in $\cal S$, i.e. by the statistically significant $l$-tuples of symbols. This is shown in Fig. 1(e) for the case of $l=2$ and $3$. In order to assess the significance of a $l$-tuple $\boldsymbol{s}$, we evaluate the expected probability $p_{\text{exp}}(\boldsymbol{s})$ from the observed occurrences of tuples of shorter length:
$p_{\text{exp}}(\boldsymbol{s}=s_{1},s_{2})$ $=p_{\text{obs}}(s_{1})p_{\text{obs}}(s_{2})$ (1)
@@
$p_{\text{exp}}(\boldsymbol{s}=s_{1},\dots,s_{l})$ $=\frac{p_{\text{obs}}(s_{1},\dots,s_{l-1})p_{\text{obs}}(s_{2},\dots,s_{l})}{p_{\text{obs}}(s_{2},\dots,s_{l-1})}$ (2)
for $l>2$. The observed probability of $\boldsymbol{s}$ in the equations above is computed as:
$p_{\text{obs}}(\boldsymbol{s})=\frac{n_{\text{obs}}(\boldsymbol{s})}{\sum_{s\in\mathcal{T}^{l}}n_{\text{obs}}(\boldsymbol{s})}=\frac{n_{\text{obs}}(\boldsymbol{s})}{S-l+1}$ (3)
where $n_{\text{obs}}(\boldsymbol{s})$ counts the number of occurrences of $\boldsymbol{s}$ in $\mathcal{D}^{l}$ and therefore satisfies $\sum_{s\in\mathcal{T}^{l}}n_{\text{obs}}(\boldsymbol{s})=S-l+1$. In this way, when evaluating $p_{\text{exp}}(\boldsymbol{s})$, we take into account the effect of lower-order correlations coming from tuples of length smaller than $l$ [43]. If we are interested in the case of unordered tuples, we can simply sum over $S_{p}(\boldsymbol{s})$, the set of unique permutations $\boldsymbol{\sigma}$ of the ordered tuple $\boldsymbol{s}=s_{1},\dots,s_{l}$, obtaining: $p_{\text{exp}}^{\text{und}}(\boldsymbol{s})=\sum_{\boldsymbol{\sigma}\in S_{p}(\boldsymbol{s})}p_{\text{exp}}(\boldsymbol{\sigma})$ and $n_{\text{obs}}^{\text{und}}(\boldsymbol{s})=\sum_{\boldsymbol{\sigma}\in S_{p}(\boldsymbol{s})}n_{\text{obs}}(\boldsymbol{\sigma})$, where $p_{\text{exp}}(\boldsymbol{\sigma})$ and $n_{\text{obs}}(\boldsymbol{\sigma})$ are respectively the expected probability and the observed occurrences of the permutation $\boldsymbol{\sigma}$ for the ordered case.
We then use a Bayesian approach to reject or accept the null hypothesis that each $l$-tuple $\boldsymbol{s}_{i}\in\mathcal{T}^{l}$ where $i\in 1,\dots,N^{l}$ appears with probability $p_{\text{exp}}(\boldsymbol{s}_{i})$. Let us define the vector of the observed count of $l$-tuples in $\mathcal{S}$: $\mathbf{n}^{l}_{\text{obs}}=\left[n_{\text{obs}}(\boldsymbol{s}_{1}),\dots,n_{\text{obs}}(\boldsymbol{s}_{N^{l}})\right]$, where $n_{\text{obs}}(\boldsymbol{s}_{i})$ is the number of times each unique tuple $\boldsymbol{s}_{i}\in\mathcal{T}^{l}$ is repeated in $\mathcal{D}^{l}$. We assume that each $l$-tuple represents the outcome of a random experiment over Sl+1S-l+1 independent trials. That is, the outcome of each trial can be tuple $\boldsymbol{s}_{i}\in\mathcal{T}^{l}$ with probability $p(\boldsymbol{s}_{i})$, such that $\sum_{i}p(\boldsymbol{s}_{i})=1$. Under this assumption, the likelihood of $\mathbf{n}^{l}_{\text{obs}}=\left[n_{\text{obs}}(\boldsymbol{s}_{1}),...,n_{\text{obs}}(\boldsymbol{s}_{N^{l}})\right]$ is a multinomial distribution:
$\mathcal{L}(\mathbf{n}_{\text{obs}}^{l}\mid\mathbf{p}^{l})=\frac{(S-l+1)!}{\prod_{i=1}^{N^{l}}n_{\text{obs}}(\boldsymbol{s}_{i})!}\prod_{i=1}^{N^{l}}p(\boldsymbol{s}_{i})^{n_{\text{obs}}(\boldsymbol{s}_{i})}$ (4)
where $\mathbf{p}^{l}=\left[p(\boldsymbol{s}_{1}),\dots,p(\boldsymbol{s}_{N^{l}})\right]$ is the vector of the occurrence probability of each tuple. For the prior $\Pi(\mathbf{p}^{l})$, as commonly done, we take the conjugate distribution of the multinomial distribution, that is the Dirichlet distribution:
@@
where $\alpha_{i}^{l}$ is the concentration parameter of the tuple $\boldsymbol{s}_{i}$, which can be interpreted as prior counts or pseudo-counts, and $\Gamma(\cdot)$ is the Gamma function. We adopt an empirical Bayes approach, where part of the information of the data is used to compute the prior. Namely, to compute the concentration parameter of an $l$-tuple, we use the expected probability in Eqs. (1),(2), functions of the observed probability of l1l-1 and l2l-2-tuples: $\alpha_{i}^{l}=n_{\text{exp}}(\boldsymbol{s}_{i})+\epsilon$. Here, $n_{\text{exp}}(\boldsymbol{s}_{i})=p_{\text{exp}}(\boldsymbol{s}_{i})(S-l+1)$ is the expected number of occurrences of the $l$-tuple $\boldsymbol{s}_{i}$ in the sequence, and $\epsilon\geq 0$ is a regularization parameter (a minimum number of pseudo-counts) that ensures that the prior is not too confident, in particular for very small $n_{\text{exp}}(\boldsymbol{s}_{i})$, and allows for Bayesian updating. For our results we choose $\epsilon=1$. We then use the likelihood to update the prior obtaining the posterior $P(p\mid\text{Data})$ which, given that the multinomial and the Dirichlet distributions are conjugate, is also a Dirichlet distribution:
$P(\mathbf{p}\mid\text{Data})=\text{Dirichlet}($ $\alpha_{1}^{l}+n_{\text{obs}}(\boldsymbol{s}_{1}),\alpha_{2}^{l}+n_{\text{obs}}(\boldsymbol{s}_{2}),$
$\dots,\alpha_{N^{l}}^{l}+n_{\text{obs}}(\boldsymbol{s}_{N^{l}}))$ (6)
@@
To assess the significance of the $l$-tuple $\boldsymbol{s}_{i}$, we compute the Jensen-Shannon distance $d_{\text{JS}}(\boldsymbol{s}_{i})$ between the marginal distribution of the prior and that of the posterior[44, 45, 46]. Notice that, the marginal of the Dirichlet distribution $(p(\boldsymbol{s}_{1}),\dots,p(\boldsymbol{s}_{N}))\sim\text{Dirichlet}(\alpha_{1},\alpha_{2},\dots,\alpha_{N})$ is the Beta distribution:
$p(\boldsymbol{s}_{i})\sim\text{Beta}(\alpha_{i},\alpha_{0}-\alpha_{i})$ (7)
@@
where $\alpha_{0}=\sum_{i}\alpha_{i}$. If $d_{\text{JS}}(\boldsymbol{s}_{i})$ is large, the null hypothesis that the probability of observing $\boldsymbol{s}_{i}$ can be estimated from the probabilities of the shorter length tuples in $\boldsymbol{s}_{i}$ must be rejected. Hence, the $l$-tuple shows pure correlations of order $l$, is considered a $l$-motif and, as such, is an $l$-hyperedge (a hyperedge with $l$ nodes) of the hypergraph $\mathcal{H}(\mathcal{N},\mathcal{E})$ of higher-order relations in the original time series. Our method naturally allows to associate weights in $[0,1]$ to the hyperedges, as the JS distance ranges between 0 and $1$. Being $d_{\text{JS}}$ a metric, it does not take into account for the relative position of the prior and posterior distributions. We therefore associate to the hyperedge $\boldsymbol{s}_{i}$ a weight equal to $d_{\text{JS}}(\boldsymbol{s}_{i})$ only when the posterior is concentrated around larger values than the prior. That is, we consider as potentially significant only those motifs for which the mean of the prior is smaller than the mean of posterior, but a similar analysis could be done also for under-represented motifs. An unweighted version of the hypergraph $\mathcal{H}$ can then be obtained by retaining only $l$-hyperedges whose weights are larger than a given threshold $d_{\text{JS}}^{\text{thr}}(l)$ (see SM for a discussion on how to choose such a threshold). In this way, the Jensen-Shannon distance between the prior and a posterior distributions of each tuple becomes a score of statistical significance; hereinafter we will indicate it as the BJS-score (Bayesian-Jensen-Shannon).
@@
Fig. 2 shows the Receiver Operating Characteristic (ROC) and Precision-Recall (PR) curves illustrating the performance of the BJS-score and the $z$-score in detecting motifs, for noise distributed according to Zipf’s law with exponent $\gamma=1$. See the Supplemental Material (SM) for similar results for other values of $\gamma$. From the left to the right, the curve points correspond to values of decreasing inference thresholds (from $1$ to 0 for the BJS, from $\infty$ to 0 for the z-score). The quicker the ROC curve rises to the upper left corner, the better the inference. On the other hand, the slower the Precision-Recall curve decreases to the bottom right corner, the better. We are interested in particular in having a high precision with few false positive. Our dataset is unbalanced, with many true negative (and the number of true negatives increases as the length of the sequence, i.e. the $r_{\text{ns}}$ increases) therefore the Precision-Recall curve is more informative. The results show that our method outperforms the $z$-score in all scenarios and is capable of detecting motifs even in extreme cases with a noise-to-signal ratio of $r_{ns}=100$. Fig. 2 shows the ROC and PR curves for the combined recognition of $2$- and $3$-motifs. For a separate analysis of $2$- and $3$-motifs, see the Supplemental Material (SM). Moreover, we found that for all the values of $\gamma$ explored, the optimal threshold for the BJS-score, corresponding to the maximum value of F1 score, is $\text{BJS}^{\text{thr}}(l)\simeq[0.5,0.7]$ for both $l=2$ (22-tuples) and $l=3$ (33-tuples). $\text{BJS}^{\text{thr}}\geq 0.5$ has also an intuitive interpretation; as the Jensen-Shannon distance takes value in $[0,1]$, where 0 corresponds to identical distributions and $1$ to completely different ones. Hence a distance greater than the middle point $0.5$ can be intuitively seen as two distributions that are more different than similar. In the next session on applications to real-world datasets, we will adopt a single threshold $\text{BJS}^{\text{thr}}=0.6$ for both $l=2$ and $l=3$.
@@
For the application to the brain, we studied the spiking activity at the level of individual neurons in the mouse during visual tasks. The multivariate time-series were obtained from the electrophysiological recordings in the Neuropixel datasets of the Allen Brain Observatory [48]. We set $x_{i}(t)=1$ to indicate that an action potential occurred in the neuron $i$ at time $t$. We considered two different spatial scales, one that involves individual neurons ($N=346$), and the other that aggregates neurons in the same functional area ($N=15$, see SM). Higher-order motifs were present in both situations. However, only at the macro-scale of functional brain areas the number $E_{3}$ of $3$-motifs exceeds the number $E_{2}$ of pairwise interactions (more than $70\%$ of the total number of hyperedges). These results are in agreement with a recent study of large-scale electroencephalographic (EEG) brain signals [34], where a similar percentage of brain dynamics was considered HO. The fact that the presence of 3 motifs significantly increases when moving from the microscale of the single neurons to the macroscale of the functional brain areas, suggests a non-negligible spatial effect that should be taken into account in future studies.